Oriole, Inc., has revenue of $471,000, costs of $329,700, and a tax rate of 31 percent. If the firm pays out 43 percent of its earnings as dividends every year, how much earnings are retained and what is the firm’s retention ratio? (Round answers to 0 decimal places e.g. 1,575.)

Retained earnings $ enter the retained earnings in dollars rounded to 0 decimal places
Retention ratio enter the retention ratio in percentages rounded to 0 decimal places %
Answer #1
Revenue $ 471,000
Less: Costs $ (329,700)
Profit before tax $ 141,300
Less: Tax $141,300 × 31% $                   (43,803)
Profit after tax $                      97,497
Less: Dividend $97,497 × 43% $                   (41,924)
Retained earnings $                      55,573
Retention ratio 57%
  • Retained earnings are the profits retained after the distribution of dividends to the investors.
  • Retention ratio shows the percentage of net income (also called profit after tax) of the company that is retained in the business.
  • Retention ratio is the opposite of the payout ratio.
  • In this question payout ratio is 43%
  • The formula for retention ratio:
    • Retention ratio = (Net income - Dividends) ÷ Net income
      • (97,497 - 41,924) ÷ 97,924 = 0.57 or 57%
    • Retention ratio = Retained earnings for the year ÷ Net income
      • 55,573 ÷ 97,497 = 0.57 or 57%
    • Retention ratio = 1 - Payout ratio
      • 1 - 43% = 0.57 or 57%